Please help me.
$2000 deposited into 2 banks. bank 1 pays 6% interest and bank 2 pays 5%. if $117 of interest was earned, how much was deposited in each bank?
I can't figure this out. Help me

Answer by macston(5194)

S=amount in 6% account; F=amount in 5% account
S+F=$2000
S=$2000-F
0.06S+0.05F=$117 Substitute for S
0.06($2000-F)+0.05F=$117
$120-0.06F+0.05F=117 Subtract $117 from each side.
$3-0.01F=0 Add 0.01F to each side.
$3=0.01F Divide each side by 0.01.
$300=F ANSWER 1 There was $300 in the 5% account.
S=$2000-F=$2000-$300=$1700 ANSWER 2:There was $1700 in the 6% account.
CHECK:
0.06S+0.05F=$117
0.06($1700)+0.05($300)=$117
$102+$15=$117
$117=$117